OpenStudy (anonymous):

Which statement is true about the sum √50+√18?

OpenStudy (anonymous):

where?

OpenStudy (anonymous):

A.) It is rational and equal to 2. B.) It is rational and equal to 8. C.) It is irrational and equal to 2√2. D.) It is irrational and equal to 8√2.

OpenStudy (anonymous):

d :)

OpenStudy (anonymous):

with this you might want to use calculator :)

OpenStudy (anonymous):

\[\sqrt{50}=5\sqrt{2}\]

OpenStudy (anonymous):

\[\sqrt{18}=3\sqrt{2}\]

OpenStudy (anonymous):

I checked, thank you for your help

OpenStudy (anonymous):

ur welcome